### v1

# LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation

Abstract
The global COVID-19 pandemic underscored major deficiencies in traditional healthcare systems, hastening the advancement of online medical services, especially in medical triage and consultation. However, existing studies face two main challenges. First, the scarcity of large-scale, publicly available, domain-specific medical datasets due to privacy concerns, with current datasets being small and limited to a few diseases, limiting the effectiveness of triage methods based on Pre-trained Language Models (PLMs). Second, existing methods lack medical knowledge and struggle to accurately understand professional terms and expressions in patient-doctor consultations. To overcome these obstacles, we construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC), comprising a Coarse-grained Triage dataset with 439,630 samples, a Fine-grained Diagnosis dataset with 199,600 samples, and a Medical Consultation dataset with 472,418 items, thereby addressing the data shortage in this field. Moreover, we further propose a novel triage system that combines BERT-based supervised learning with prompt learning, as well as a GPT-based medical consultation model using reinforcement learning. To enhance domain knowledge acquisition, we pre-trained PLMs using our self-constructed background corpus. Experimental results on the LCMDC demonstrate the efficacy of our proposed systems. Our corpora and codes are available at: Corpora and Codes
LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation
Xinyuan Wang  and Haozhou Li  and Dingfang Zheng  and Qinke Peng Xi’an Jiaotong University {wxy0713, lihaozhou1126, zdf0921}@stu.xjtu.edu.cn qkpeng@mail.xjtu.edu.cn
1 Introduction
The outbreak and spread of COVID-19 have placed immense pressure on global healthcare systems, exposing numerous shortcomings in traditional healthcare services Sun et al. (2021). In response to the urgent need for pandemic control, online medical consultations and tele-medicine have become crucial for maintaining healthcare services Jnr (2020). Between July and December 2022, 58.5% of U.S. adults used the Internet to seek health or medical information, 41.5% communicated with a doctor or doctor’s office, and 46.1% checked medical test resultsWang and Cohen (2023). According to the China Internet Network Information Center (CNNIC), as of December 2023, the number of Internet medical users in China reached 414 million. This trend reflects the growing reliance on digital platforms for healthcare.
The rise of online healthcare platforms has made medical triage and intelligent Q&A key research areas, helping reduce the burden on healthcare systems by providing remote medical advice and reducing unnecessary hospital visits. These technologies enhance resource efficiency, minimize cross-infection risks, and improve access to healthcare services, especially in remote areas.
However, most existing studies face two main challenges. Firstly, due to data privacy concerns, obtaining large-scale public medical triage and consultation datasets is difficult. Although some studies have collected corpora for automatic diagnosis Wei et al. (2018); Lin et al. (2019); Xu et al. (2019) and dialogue generation Liu et al. (2022), these datasets are limited to a few specific diseases and are small in size Zeng et al. (2020), making them inadequate for training intelligent medical diagnosis models based on pre-trained language models. Secondly, patient-doctor consultations usually utilize distinct expressions and professional terms, making it difficult for existing methods that lack domain-specific knowledge to comprehend medically pertinent information Li et al. (2024).
To address the limitations of existing datasets, we have constructed the Large-scale Chinese Medical Dialogue Corpora (LCMDC) by extracting patient-doctor conversations from the "Quick Doctor" website  [^1] https://www.120ask.com/. Our work addresses two key needs in online healthcare: task-oriented medical triage and open-ended consultation systems. We developed a coarse-grained triage dataset with over 430,000 patient-doctor dialogues across 14 departments and a fine-grained diagnosis dataset with 200,000 entries covering 120 diseases. For medical consultation, we compiled over 470,000 question-answer pairs. Additionally, we collected disease information datasets to enrich Pre-trained Language Models (PLMs) with medical knowledge for pre-training.
Nowadays, natural language processing techniques are increasingly effective in intelligent medical triage and dialogue generation tasksSun et al. (2017). Specifically, neural network-based text classification algorithms with attention mechanisms have been successfully applied to medical triage and diagnosisWei et al. (2018); Lin et al. (2019); Xu et al. (2019). However, existing studies often concentrate on a limited number of diseases and struggle with imbalanced label distributions in many-class medical diagnoses Zeng et al. (2020). Although some pre-trained language models, such as BERT and ERNIE, have been employed in automatic triage systems Wang et al. (2021); Wang and Feng (2022), they lack pre-training on medical data, hindering their ability to grasp domain-specific knowledge Li et al. (2024). To overcome these limitations, we propose a novel triage system that combines PLM-based supervised learning with prompt learning methods. We initially pre-train BERT using a self-built medical background corpus, then employ medical BERT, BiLSTM, and Dendrite Network Liu and Wang (2021) to represent the input questions for classification in our Triage datasets. Additionally, we incorporate prompt learning to mitigate the many-label imbalance issue, enhancing prediction accuracy for rare diseases.
In the medical text generation task, sequence-to-sequence models, including LSTM-based Seq2Seq methods Liu et al. (2022) and transformer-based PLMs like BERT-GPT Zeng et al. (2020), have been widely used, and achieved promising results. However, traditional LSTM models, which generate output solely based on input sequences, struggle with insufficient information utilization due to their gating mechanisms. Similarly, conventional PLMs often fail to effectively comprehend medical information within patient-doctor conversations. Additionally, these methods typically lack supplementary information for enhancement, resulting in poor performance in generating patient-doctor dialogues. To address these challenges, we re-train GPT-2 using our collected medical background corpus to internalize domain-specific knowledge. We also construct a domain knowledge graph and extract relevant content about each user’s question as supplementary input information to enhance the model. Finally, we train our model using our Medical Consultation Dataset and fine-tune it with a reinforcement learning method to simulate human doctors in providing responses.
The contributions of this paper are three-fold:
  • Collected and integrated over millions of samples to form three Large-scale Chinese Medical Corpus. To support medical domain language models and lay the foundation for automatic triage and medical consulting systems.
  • Proposed an information fusion-based classification algorithm and a prompt learning classification algorithm to improve intelligent triage accuracy, especially for rare diseases, achieving a 5% improvement in accuracy.
  • Developed a dialogue model construction framework including knowledge incorporation and input supplementation, improving performance across multiple evaluation metrics.
2 Chinese Medical Corpus
To address the scarcity of large-scale corpora and the constraints of current datasets, we collected millions of doctor-patient conversations to construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC). This collection is designed to enhance medical triage and dialogue generation and includes three components: 1) a Coarse-grained Triage dataset, 2) a Fine-grained Diagnosis dataset, and 3) a Medical Consultation dataset.
2.1 Data Collection
To make the model more adaptable to conversational and ambiguous medical descriptions, we claw the original data from the online medical consultation platform named "Quick Doctor". Reasons are shown in Appendix B.1.
(a) Age Distribution
(b) Gender Distribution
Figure 1: Age and Gender Distribution in our Corpus.
We collect millions of entries containing doctor responses and conduct statistical analysis and visualization based on different labels to better understand the original data. Our analysis, illustrated in Figure 1, explores the gender and age of users. The results show that female users significantly outnumber male users. Predominantly, young individuals aged 10-40, who are more likely to incorporate the Internet into their daily routines, made the inquiries.
2.2 Intelligent Triage Dataset
For automatic medical triage and diagnosis, we construct two datasets of different granularities for experimental analysis. The triage process is conceptualized as a text classification task, where patient consultations serve as input, and department classifications function as labels.
We organize the original corpus into two datasets after data preprocessing: 1) a Coarse-grained Triage dataset with primary department labels for higher accuracy in quickly identifying potential conditions and directing patients to the appropriate department, and 2) a Fine-grained Diagnosis dataset with tertiary department labels for more precise guidance, thereby helping users access detailed medical knowledge and improving query efficiency. The Coarse-grained dataset has over 430,000 entries with 14 categories, while the Fine-grained dataset has over 190,000 entries with 120 categories, as shown in Table 1. Based on these datasets, we further propose an intelligent triage algorithm based on supervised learning and prompt-based learning and then evaluated its effectiveness.
Table 1: Statistic of Intelligent Triage Datasets
Dataset Name Coarse-grained Fine-grained
Number of Data 439,630 199,600
Training Set 373,686 169,660
Test Set 65,944 29,940
Average Sentence Length 57.77 58.30
Number of Categories 14 120
2.3 Medical Consultation Dataset
For the medical dialogue generation, we extract the Question-Answer sample from the original corpus to create the Medical Consultation Dataset, which can offer several advantages: 1) Major medical Q&A portals have a large user base with comprehensive questions, providing a wide range of input data for the model to learn from; 2) High-quality doctors registered on these sites provide accurate responses, providing high-quality domain knowledge for training a professional model; 3) The large data volume allows extensive training to improve model performance from various perspectives.
Table 2: Statistic of Medical Consulting Dataset
Dataset Item Quantity
Q&A Data Total Number of Q&A 472,418
Training Set Number of Entries 467,418
Average Question Length 57.86
Average Answer Length 92.57
Test Set Number of Entries 5,000
Average Question Length 57.88
Average Answer Length 92.54
Specifically, each sample in this dataset comprises a patient consultation and a doctor’s response, serving respectively as the model’s input and prediction target. We conducted data cleaning to remove entries that lacked either questions or answers and eliminated entries where the questions and responses were shorter than 10 characters, thereby enriching the training data with more substantive content. After data preprocessing, the Medical Consultation Dataset contains a total of 472,418 samples, with an average question length of 57.86 characters and an average response length of 92.57 characters. The statistical results of the dataset are presented in Table 2.
3 Intelligent Triage System
We propose two text classification-based methods to evaluate intelligent triage datasets, using coarse-grained and fine-grained data to provide varying levels of consultation suggestions. We re-train language models on medical text data to improve their performance on Chinese medical texts and develop a supervised text classification model based on these domain-adapted models. For categories with limited data, a prompt-based text classification method leverages pre-trained language models’ knowledge, enhancing performance on small datasets.
3.1 Problem Definition
The intelligent medical triage task is modeled as a text classification problem, processing user queries to determine the appropriate department for consultation.
$p\left\{{c_{1},~{}c_{2},\ldots,c_{k_{l}}}\right\}=TextCls\left({t_{1},t_{2},\ldots,t_{k_{w}}}\right),$ (1)
where $p\left\{{c_{1},~{}c_{2},\ldots,c_{k_{l}}}\right\}$ means the probability on each label.
3.2 Supervised Learning Triage System
First, we propose a traditional supervised learning-based text classification method. The proposed network framework is shown in Figure 2. The input is the user query text, and the output is the classification distribution of the text. First, we use a pre-trained language model (e.g., BERT) to obtain the CLS vector of the text, which is the classification vector inherent in BERT. Simultaneously, we extract the token representation vectors encoded by BERT and use a bidirectional LSTM network to capture temporal information. The CLS vector and the output vectors from the LSTM layer are then fused and processed using a Dendritic Network for information integration. Finally, a Single-Layer Perceptron and a Softmax function are used to obtain the classification distribution.
Figure 2: The Proposed Supervised Learning Classification Method.
The model consists of four steps: The first step is domain-specific knowledge learning based on pre-training. To make the pre-trained language model used in this model more suitable for the medical domain, we further pre-trained the general pre-trained language model using the collected medical encyclopedia data and medical consulting data. Self-supervised learning was performed using the MLM pre-training task, enabling the model to better learn medical knowledge and colloquial expressions. The pre-training process employed the same parameters as BERT, selecting 15% of the total words as masked tokens. The training objective of the MLM task is to minimize the following negative log-likelihood loss function:
$L=-{\sum\limits_{\hat{x}\in m{(x)}}{\log{P\left(\left.\hat{x}\right|\right.}}}\left.X_{\backslash m{(x)}}\right)$ (2)
Based on the pre-trained language model with medical text supplementation, we represent the text by obtaining the classification vector inherent to the pre-trained language model and the text token representation vector encoded by the pre-trained language model.
$\mathbf{X}_{i}^{emb}=BERTemb\left(T_{i}\right)$ (3)
The embedding vectors are then computed by the multi-head self-attention mechanism in BERT to get the representation vectors and the CLS vector.
$\mathbf{S}_{i}^{l,j}=softmax\left(\frac{\mathbf{Q}_{i}^{l,j}*{\mathbf{K}_{i}^{l,j}}^{T}}{\sqrt{k_{h}}}\right)*\mathbf{V}_{i}^{l,j}~{},~{}$ (4)
where $\mathbf{S}_{i}^{l,j}\in\mathbb{R}^{k_{m}\times k_{h}}$ are representations with attention.
${\mathbf{O}_{i}^{l}}_{(1)}=LayerNorm\left({\mathbf{X}_{i}^{l-1}+{\mathbf{S}_{i}^{l}}_{(2)}}\right)~{},$ (5)
where ${}{\mathbf{O}_{i}^{l}}_{(1)}\in\mathbb{R}^{k_{m}\times k_{h}}$ are updated representations.
Next, the temporal information is extracted using a bidirectional LSTM network
${\overset{\rightarrow}{\mathbf{H}}}_{i}$ $=\overset{\rightarrow}{LSTM}\left(\mathbf{P}_{i}\right)$ (6)
${\overset{\leftarrow}{\mathbf{H}}}_{i}$ $=\overset{\leftarrow}{LSTM}\left(\mathbf{P}_{i}\right)$
$\mathbf{H}_{i}$ $=\left[{~{}{\overset{\rightarrow}{\mathbf{H}}}_{i},~{}{\overset{\leftarrow}{\mathbf{H}}}_{i}}\right],~{}\mathbf{H}_{i}\in\mathbb{R}^{2\bullet k_{h}}$
After obtaining the representation $\mathbf{H}_{i}$ summarized by the LSTM network, we fuse it with the BERT output classification vector $[\text{Cls}]_{i}$ to form a new representation vector $\mathbf{M}_{i}$:
$\mathbf{M}_{i}=[\mathbf{H}_{i};\text{Cls}_{i}]$ (7)
This vector is then input into the DD layer for processing, which reduces the probability of overfitting and improves the model’s robustness.
$\mathbf{D}_{i}=\mathbf{W}_{\mathbf{d}}*\left({\mathbf{M}_{i}\bullet\mathbf{M}_{i}}\right),\mathbf{W}_{\mathbf{d}}\in\mathbb{R}^{{({3\bullet k_{h}})}\times k_{h}}$ (8)
Finally, the decision is made through a dense layer:
$\mathbf{C}_{p}=Softmax\left({\mathbf{D}_{i}*\mathbf{W}_{F}+\mathbf{b}_{F}}\right),$ (9)
where $\mathbf{W}_{F}\in\mathbb{R}^{k_{h}\times k_{l}}$ and $\mathbf{b}_{F}\in\mathbb{R}^{k_{l}}$. And then the loss function used here is:
$loss=-{\sum\limits_{i=1}^{k_{l}}{c_{i}{\mathit{\log}\left(p_{i}\right)}}}$ (10)
3.3 Prompt Learning Triage System
Both granularity-level classification datasets suffer from data imbalance issues. For categories with less data, the performance of conventional text classification methods encounters certain bottlenecks. To address this, we propose a prompt-based classification method that leverages the knowledge embedded in pre-trained language models tailored to the medical domain to improve classification accuracy. In this prompt-based classification method, the label is a list of text, and the classification process involves predicting the content of this text, thus achieving classification. This approach establishes a relationship between the input sentence and the label name. It effectively utilizes the knowledge within pre-trained language models and performs well on datasets with few samples, making it suitable for classifying rare diseases with limited samples. Specifically, for each user input text, we add a prompt sentence with a masked word, shown in Figure 3.
Figure 3: The Proposed Prompt Structure for Classification.
In this way, the classification task transforms into predicting the masked words. Interestingly, this training method aligns with the MLM task in BERT’s pre-training process, allowing us to leverage the knowledge acquired during pre-training for fill-in-the-blank tasks. This accelerates the training process and is suitable for categories with fewer training samples.
4 Medical Consultation System
For the medical consultation dataset, we developed a Chinese medical dialogue system using a generative language model to simulate interactions with doctors. The system incorporates medical knowledge and supplements inputs, with user questions as input text and doctor’s responses as output text for training.
4.1 Problem Definition
We mathematically model building a consultation system as a text-generation task. Specifically, based on the input question $Q=(q_{1},q_{2},\ldots,q_{M})$, we use a knowledge graph for supplementary input, denoted as $I=Graph(Q)=(i_{1},i_{2},\ldots,i_{L})$. The text generation model produces the response sequence $A=(a_{1},a_{2},\ldots,a_{N})$ as follows:
$\mathbf{A}$ $=TextGen\left(\mathbf{Q},\mathbf{I}\right)$ (11)
$=TextGen\left(q_{1},q_{2},\ldots,q_{M},i_{1},i_{2},\ldots,i_{L}\right)$
4.2 Overview
Based on the GPT-2 model, we re-train the model using the collected medical encyclopedia text data to internalize knowledge, making it more suitable for tasks in the medical domain. Next, we construct a medical knowledge graph and use a search-and-match approach to find relevant knowledge to each question, providing external information as supplementary input text. Then, we train the model using dialogue data and finally fine-tune it using reinforcement learning.
According to the above process, we construct the domain-specific medical consultation system, as shown in Figure 4.
Figure 4: The Medical Dialogue System Framework.
4.3 Knowledge Injection Answer Generation
Knowledge injection allows the model to have rich prior knowledge, which involves re-training the large language model using the collected basic knowledge text data. After such training, the originally broad language model can focus more on the medical field, incorporating this medical knowledge to support subsequent medical consultation.
In the re-training stage, we view a sentence as a token sequence: $Sentence=\left({{token}_{1},{token}_{2},\ldots,{token}_{N}}\right)$. The objective is to maximize the following objective function:
$L\left(Sentence\right)=\sum\limits_{n=K}^{N}p\left({token}_{n}\,\middle|\,{token}_{n-K},\ldots,{token}_{n-1};\theta\right)$ (12)
where $\theta$ represents the parameters in the GPT-2 model that need to be fine-tuned, and encompass knowledge.
Knowledge injection is crucial because large models are trained on diverse datasets, making them broad but not sufficiently focused on specific domains. Internalizing knowledge helps the model concentrate more on the required domain and incorporate more domain knowledge. Next is the input supplementation process, which involves constructing a knowledge graph and searching for relevant text information based on the input text. This information, combined with the input question text, serves as new input for external knowledge supplementation. This process simulates how humans search for relevant knowledge based on the question and respond using both the question and related knowledge. The subsequent stage involves training the model with domain-specific data. This stage aims to teach the model the dialogue patterns and generate responses based on input data. This corresponds to the fine-tuning phase of the pre-trained language model, enabling the model to simulate human doctors in responding.
5 Experiments
We build the intelligent triage system and medical consultation system and test them on the collected datasets.
5.1 Experiment Setup
The experiments for the triage and consultation systems, including datasets, baselines, evaluation metrics, hyper-parameters, and environment settings, are outlined in detail in the Appendix C.
5.2 Supervised Triage System Experiment
5.2.1 Overall Performance
For the supervised learning classification method, we employ different baselines including machine learning methods, neural network methods, and pre-trained models. The results of different-grained datasets are shown in Table 3
Table 3: Experimental Results on Different Granularity Datasets
Coarse-grained Dataset Results Fine-grained Dataset Results
Classification Method Model Accuracy F1 Score Precision Recall Accuracy F1 Score Precision Recall
Machine Learning SVM 57.27% 44.27% 46.67% 42.10% 62.68% 53.83% 56.75% 51.19%
Decision Tree 61.24% 48.86% 52.47% 45.71% 59.68% 45.61% 45.49% 45.74%
Neural Network TextCNN 77.07% 54.36% 56.83% 52.10% 76.38% 61.29% 66.92% 56.54%
RCNN 71.56% 51.40% 53.76% 49.24% 73.84% 59.86% 61.10% 58.67%
LSTM 73.38% 53.02% 55.37% 50.86% 71.59% 59.80% 66.91% 54.08%
Self Attention 64.52% 52.17% 54.22% 50.27% 68.08% 60.17% 64.90% 56.08%
Capsule 62.70% 51.47% 53.01% 50.01% 65.98% 52.23% 57.88% 47.59%
Pre-trained Model BERT 79.60% 59.92% 62.51% 58.79% 80.53% 65.65% 70.64% 65.24%
RoBERTa 77.59% 55.69% 60.23% 53.98% 77.31% 55.74% 61.37% 53.46%
MacBERT 80.17% 59.68% 66.56% 59.84% 80.73% 65.23% 71.13% 64.88%
Experimental Algorithm No PT Benchmark 81.34% 63.20% 65.05% 62.76% 82.47% 72.01% 75.60% 69.78%
Our Method 82.05% 65.60% 66.54% 65.72% 83.31% 74.12% 77.31% 72.30%
According to the experimental results, the proposed supervised learning-based text classification algorithm demonstrates better classification performance compared to machine learning and deep neural network-based text classification algorithms. It shows improvements in metrics such as accuracy and F1-score, indicating the effectiveness of the proposed algorithm. Additionally, compared to commonly used pre-trained language models, the proposed classification algorithm leverages more information and effectively integrates this information, leading to enhanced classification performance.
When comparing the classification results on the coarse-grained and fine-grained datasets, it is found that the classification performance on the fine-grained dataset, which has more categories, is actually better. This issue can be attributed to the nature of the datasets themselves. Both granularity-level datasets are derived from the same batch of collected data, with the coarse-grained dataset containing over 430,000 entries, while the fine-grained dataset contains only about 190,000 entries. Since the fine-grained dataset uses tertiary department or disease labels, these labels provided by human doctors are very precise. However, the coarse-grained dataset only requires primary labels, leading to many texts in the coarse-grained dataset being ambiguously classified, making it difficult even for human doctors to provide suitably detailed labels. This results in better classification outcomes on the fine-grained dataset compared to the coarse-grained dataset.
Ablation study results are presented in Appendix D.1
5.2.2 Results of Domain Information Injection
To investigate the differences before and after domain-specific re-training of different pre-trained models, we designed comparative experiments to evaluate model performance under different conditions. We used the original BERT and MacBERT models, as well as the proposed models utilizing these pre-trained models, comparing their performance before and after re-training. The results are shown in Figure 5
(a) Coarse-grained
(b) Fine-grained
Figure 5: Results of Domain Information Injection.
With the same pre-trained models and classification network structures, re-training the language model can improve the accuracy of the classification model. This is because the re-training data is related to Chinese medical information, allowing the language model to learn the intrinsic knowledge of medical texts, resulting in better performance on tasks involving medical texts.
From this set of comparative data, it can be seen that using re-trained language models significantly improves accuracy and F1 scores. The proposed model, combined with the re-trained language models, achieves even better classification results. This indicates that re-training the language model with domain-specific data enhances its representation capability for domain-specific tasks. In this experiment, the language model re-trained with medical knowledge and Q&A text data demonstrates better representation capability for texts that integrate colloquial expressions and medical knowledge.
5.2.3 Parameters Study
We also investigated the impact of different numbers of dendritic neural network layers and LSTM layers on classification results to select the optimal algorithm model structure.
(a) DD layer
(b) LSTM layer
Figure 6: Results of DD and LSTM layer Numbers.
According to the results in Figure 6, the dendritic neural network achieves the best classification performance with 3 layers, and the LSTM network achieves optimal performance with 2 layers.
5.3 Prompting Triage System Experiment
Prompt learning quickly completes tasks using pre-trained knowledge, even with limited data. To leverage this advantage, we created small sample datasets by removing high-data categories from the coarse-grained and fine-grained datasets, resulting in 73,604 entries across nine categories and 24,940 entries across 83 categories, respectively.
Table 4: Prompt Learning Classification Results on Different Granularity Datasets
Dataset Method Model
Original BERT Re-trained BERT Supervised Learning
Coarse-grained Original 77.47% 79.39% 79.60%
Small Sample 86.36% 88.18% 87.16%
Fine-grained Original 79.30% 81.73% 83.31%
Small Sample 79.28% 80.11% 77.31%
Table 4 shows that the BERT model re-trained with collected data performs better in prompt learning-based classification, surpassing supervised learning on small datasets.
However, prompt learning has drawbacks: it requires more training time due to predicting multiple token positions, and its performance can be slightly worse than supervised learning. This is due to the need to correctly predict all token positions and handle varying label lengths with padding symbols, which can interfere with pre-trained knowledge and reduce accuracy. Additionally, prompt learning is less effective with large datasets.
5.4 Question Answer System Experiment
5.4.1 Overall Performance
First, we compared the proposed framework’s text generation model with a deep learning-based sequence-to-sequence (seq2seq) model. The seq2seq model is constructed with LSTM networks for both the Encoder and Decoder and includes information interaction methods such as attention mechanisms and information fusion representations. This model serves as a representative of deep learning methods that do not incorporate background information.
Additionally, we compared the proposed text generation model with the BART model. BART combines features of both autoencoding and autoregressive language models and is widely used in the field of text generation, especially for text summarization. In this study, we utilized the general Chinese BART model to build a dialogue system and trained it using medical dialogue data.
Table 5: Comparison of Text Generation Results
Evaluation Metric Seq2Seq Model BART Our Model
Weight F1 $\uparrow$ 0.1275 0.2070 0.2087
Weight P $\uparrow$ 0.1569 0.2215 0.2214
Weight R $\uparrow$ 0.1321 0.1998 0.2102
BLEU-1 $\uparrow$ 0.0922 0.1214 0.1211
CHRF $\uparrow$ 0.0374 0.0769 0.0781
GLEU $\uparrow$ 0.0448 0.0977 0.1029
NIST $\uparrow$ 0.5398 1.3932 1.4049
RIBES $\uparrow$ 0.0637 0.0827 0.0903
TER $\downarrow$ 1.3519 1.0489 0.9997
WMD $\uparrow$ 0.5168 0.6192 0.6461
BERT F1 $\uparrow$ 0.568 0.646 0.648
BERT P $\uparrow$ 0.576 0.649 0.652
BERT R $\uparrow$ 0.562 0.644 0.644
Information Entropy $\uparrow$ 5.4215 5.7422 5.7291
Lexical Diversity $\uparrow$ 0.018 0.149 0.188
KL Divergence – 0.7888 0.0549 0.0619
Self BLEU Bigram $\downarrow$ 0.9804 0.9875 0.9870
Self BLEU Trigram $\downarrow$ 0.9581 0.9694 0.9683
Web Bert Similarity $\uparrow$ 2.2324 2.2858 2.2960
According to Table 5, the proposed model outperforms the traditional deep learning sequence-to-sequence (seq2seq) algorithm across almost all evaluation metrics. TER (Translation Edit Rate) measures the number of edits required to change a generated sentence to match the reference sentence, indicating the proximity between the generated and real texts. Therefore, a lower TER value implies a closer match between the generated and real texts. The superior performance of the model on these evaluation metrics is attributed to the knowledge embedded in the large model and the rationality of the proposed framework.
The Self-BLEU metric measures the BLEU score of a generated text against itself. A lower Self-BLEU score indicates greater diversity in the generated text, suggesting that the model produces more varied and information-rich content. However, it was observed that the seq2seq model had a lower Self-BLEU score than the proposed model. This is because the seq2seq model generates texts with many uncommon characters, increasing the perceived richness of the generated text. However, these characters do not add actual meaningful content, as reflected in the other metrics.
Ablation study results are presented in Appendix D.2
6 Discussion and Future Work
We discuss the biases in online medical data, AI adoption in healthcare, and the development of large language models in different sectors. For details, please refer to Appendix E.
Looking ahead, future work focuses on expanding datasets to reduce bias, enhancing intelligent triage systems with background knowledge and multi-modal methods, and refining medical consultation systems through alignment between LLM outputs and human preferences. The potential for further applications of LLMs is vast, enabling more flexible and realistic solutions in various fields. For details, please refer to the Appendix F.
7 Conclusion
We collected and organized Chinese medical data, including datasets for automated medical triage and consultation, forming the foundation of our work. Based on these datasets, we developed two systems: an intelligent triage system using text classification and a medical consultation system using a generative language model. Experimental results and ablation studies validated the effectiveness of both systems, demonstrating improved performance in medical diagnosis and consultation tasks.
8 Limitation
Dataset Biases:
  • Lack of Emergency Data: The dataset lacks emergency and severe cases, limiting the model’s ability to generalize to critical or urgent medical situations.
  • Imbalanced Age Distribution: The dataset is biased towards younger users, with insufficient data on elderly populations, reducing the model’s applicability to older age groups.
  • Issues Due to Anonymity: The anonymity of online consultations often leads to more private or informal questions, which may not accurately reflect typical clinical scenarios, affecting the model’s real-world performance.
Intelligent Triage System:
  • Lack of Comparison with Human Performance: Current classification methods lack sufficient comparison with professional doctors’ diagnostic results, making assessing their reliability in practical medical applications difficult.
  • Lack of Multi-modal Research: Existing classification methods are primarily text-based and do not incorporate multi-modal data such as images or audio, limiting the understanding of comprehensive patient information.
Medical Consultation System:
  • No Multi-turn Dialogues: Generation methods are often limited to single-turn dialogues, lacking the ability to handle contextual information across multiple rounds of interaction, which is crucial in real consultations.
  • No Role-playing or Simulation: There is a lack of role-playing or simulation of real medical scenarios in the generated dialogues, resulting in responses that lack authenticity and practical relevance to clinical settings.
9 Ethical Issues
Our study does not present any ethical issues that would require formal ethics approval, as it does not involve human subjects, sensitive personal data, or any other ethical considerations that would necessitate formal ethics oversight. First, the anonymized patient-doctor dialogues were legally and compliantly obtained from publicly available sources, specifically from an online medical consultation platform. Second, users have agreed to the platform’s terms of service, which include being informed that these anonymized consultation dialogues may be made public, so the data has been made available to the public with the informed consent of the users.
Appendix A Related Work
A.1 Dataset
CBLUE Zhang et al. (2022) is a comprehensive Chinese biomedical language understanding benchmark designed to advance research in biomedical NLP for the Chinese language. MedDG Liu et al. (2020) is a large-scale dataset aimed at developing conversational agents for primary clinical advice, especially valuable during the COVID-19 pandemic. MedDialog Zeng et al. (2020) is a comprehensive dataset for training and evaluating medical dialogue systems, enhancing remote healthcare services. MedQuAD Ben Abacha and Demner-Fushman (2019) is a significant resource for developing AI models in medical question answering, providing a robust foundation for understanding and responding to medical inquiries. There are some other datasets like event recognition Deng et al. (2022).
A.2 Triage System
IntelTriage Billis et al. (2019) is an intelligent triage system for Emergency Departments in Greece, dynamically prioritizing patients and monitoring their vital signs using wearable biosensors. The Babylon AI-powered Triage and Diagnostic System Razzaki et al. (2018) demonstrated diagnostic accuracy comparable to human doctors and provided generally safer triage recommendations. MESTRIMAN Sierra et al. (1993) is an expert system designed to support medical assistance during catastrophes by coordinating resources and enhancing disaster management. The SADE model Li et al. (2024) addresses challenges in medical triage and diagnosis by using dual encoders to represent patient consultations and doctor diagnoses.
A.3 Question Answer System
This research Abdallah et al. (2020) automates the generation of qualified medical answers using an end-to-end RNN and encoder-decoder model trained on data from online health services. Another system Jiang et al. (2021) integrates medical knowledge, knowledge graphs, and QA systems, using crawler technology and rule-based algorithms for classification and querying. An end-to-end character-level multi-scale CNN framework Zhang et al. (2017) is introduced for Chinese medical question-answer matching. Med-PaLM 2 Singhal et al. (2023) leverages large language model improvements and medical domain finetuning to enhance performance on medical QA datasets.
Appendix B Dataset Details
B.1 Reasons for Data Source
Here are several reasons for choosing "Quick Doctor" website:
  • Apart from sensitive information, the patient consultations and doctor responses are public, reducing the difficulty of data collection.
  • The portal boasts an extensive dataset, containing over 100 million entries, which adequately meets our requirements.
  • The portal provides comprehensive information, including user characteristics like gender and age, and detailed classifications of medical departments across four levels, facilitating robust data statistics and medical department recommendations.
Appendix C Experiment Setup Details
C.1 Datasets
In Section 2, we collect medical corpus and formulate 2 datasets for the triage system and consultation system respectively. In the intelligent triage dataset, the inputs are the users’ questions, and the outputs are two different grained labels, indicating high-level and detailed triage.
C.2 Baselines
For the triage system, we employ SVM Cortes and Vapnik (1995) and decision tree Quinlan (1986) as the mechanism learning baselines, while TextCNN Kim (2014), RCNN Lai et al. (2015), LSTM Hochreiter and Schmidhuber (1997), Self-Attention Vaswani et al. (2017), and Capsule Network Sabour et al. (2017) are used as neural network baselines. Then We also utilize pre-trained language models including BERT, RoBERTa Liu et al. (2019), and MacBERT Cui et al. (2020). For the consultation system, we employ the Seq2Seq method and BART Lewis et al. (2020) as baselines.
C.3 Evaluation Metrics
For the triage system, we utilize classification metrics, including accuracy, precision, recall, and F-1 score. For the consultation system, we employ various evaluations, including n-gram overlap measures such as BLEU Papineni et al. (2002), NIST Doddington (2002), RIBES Isozaki et al. (2010), CHRF Popović (2015), and GLEU Mutton et al. (2007); distance-based measures like TER Snover et al. (2006); vector similarity measures including WMD Kusner et al. (2015); n-gram diversity measures such as Self BLEU; sentence semantic similarity measures including BERT score Zhang et al. (2019) and Web Bert Similarity; and other metrics such as Lexical Diversity, and KL Divergence.
C.4 Hyper Parameters
For continuous pre-training of BERT and MacBERT, we set the max input length as 128 and mask 15% of tokens in the MLM task. For the supervised training of the triage system, we set the embedding dimension to 768, and the LSTM hidden size to 1024. We train 50 epochs. The learning rate for the pre-trained language model is 5e-5, while that for other parts of the whole framework is 2e-4. For the prompt learning method, we train 20 epochs with a learning rate of 2e-5. For the consultation system, we use GPT-2 as the base model. We set the token embedding dimension to 768, and the vocabulary length of natural language tokens to 50257. In the pre-training stage, we trained 10 epochs using background knowledge. In the fine-tuning stage, we trained 20 epochs using a learning rate of 2.6e-5.
C.5 Experimental Environment
All experiments were conducted on Ubuntu 22.04.3 LTS OS, with the framework of Python 3.11.5 and PyTorch 2.0.1. All data are computed on 4 NVIDIA GeForce RTX 2080 Ti GPUs, each featuring 12 GB of memory with CUDA version 12.2.
Appendix D Ablation Study Details
D.1 Supervised Triage System Experiment
To verify the necessity of each component in the supervised triage system , we conducted ablation experiments by removing each component.
(a) Coarse-grained
(b) Fine-grained
Figure 7: Ablation Results of Supervised Classification Method.
When any part of the proposed classification algorithm is removed or replaced with another structure, the classification accuracy of the new model decreases, as shown in Figure 7. This indirectly highlights the importance of each component of the proposed model.
D.2 Medical Consultation System Experiment
Next, we conducted ablation experiments to determine the importance of each component of the medical consultation system. Three comparative models were set up: a model without knowledge internalization, a model without input supplementation, and a model without both components.
Table 6: Results of Text Generation Ablation Experiments
Evaluation Metric Ours w/o Knowledge w/o Input Neither
Weight F1 $\uparrow$ 0.2087 0.2012 0.2005 0.1998
Weight P $\uparrow$ 0.2214 0.2108 0.2148 0.2111
Weight R $\uparrow$ 0.2102 0.2055 0.2016 0.2014
BLEU-1 $\uparrow$ 0.1211 0.1132 0.1181 0.1176
CHRF $\uparrow$ 0.0781 0.0745 0.0749 0.0736
GLEU $\uparrow$ 0.1029 0.0998 0.0989 0.0974
NIST $\uparrow$ 1.4049 1.3794 1.3755 1.3736
RIBES $\uparrow$ 0.0903 0.0850 0.0881 0.8237
TER $\downarrow$ 0.9997 1.0196 1.0215 1.0588
WMD $\uparrow$ 0.6461 0.6343 0.6351 0.6332
BERT F1 $\uparrow$ 0.648 0.647 0.645 0.644
BERT P $\uparrow$ 0.652 0.652 0.651 0.649
BERT R $\uparrow$ 0.644 0.643 0.641 0.641
Information Density $\uparrow$ 5.7291 5.6773 5.6226 5.6176
Lexical Diversity $\uparrow$ 0.188 0.168 0.155 0.156
KL Divergence – 0.0619 0.0784 0.0973 0.0719
Self BLEU Bigram $\downarrow$ 0.9870 0.9888 0.9896 0.9898
Self BLEU Trigram $\downarrow$ 0.9683 0.9696 0.9716 0.9729
Web Bert Similarity $\uparrow$ 2.2960 2.2855 2.2819 2.2633
The results in Table 6 demonstrate that each step in the proposed framework is crucial. Removing any step results in a decline in the model’s performance across various evaluation metrics, highlighting the superiority and rationality of the proposed structure.
The knowledge internalization step enables the generative pre-trained language model used for constructing the Q&A model to learn more medical knowledge. This is similar to how human doctors leverage their background knowledge to provide answers. Hence, the knowledge internalization step enriches the Q&A model with medical prior knowledge.
The input supplementation step simulates the process of human doctors searching for knowledge when encountering specific cases. By targeted acquisition of prior knowledge, combined with the powerful information retrieval capabilities of the knowledge graph, the model can generate more accurate response sentences. This process also contributes to the interpretability of the model construction process.
Appendix E Discussion
E.1 Data Issues
During data collection, we found certain biases in the data gathered from online platforms. Firstly, acute and severe cases are not typically presented online, as patients in these situations usually go directly to the hospital. Secondly, the internet acts as a filter, with the online population differing from the general population. This results in a majority of younger users seeking online consultations, leading to a lack of data on elderly conditions. Lastly, due to the anonymity of the internet, the nature of questions asked online can differ from those asked offline, often including more private issues.
Medical issues require maximum caution, and licensed doctors undergo extensive training to earn their qualifications due to the severe consequences of medical errors, which can endanger lives. Therefore, using current AI, which often functions as a black-box system, for diagnosis and treatment remains an ethical and legal grey area. At present, AI can only serve as a supplementary tool. Similar to autonomous driving, where accidents pose challenges in accountability, AI in healthcare must develop gradually. Automation systems must advance in a phased manner.
E.2 Discussion on Current Large Models
With the rise of ChatGPT, pre-trained language models seem to have deviated from their original purpose but have unified various tasks. Previously proposed task classifications are now outdated. Since the introduction of the T5 model, all-natural language processing tasks can be modeled as text generation tasks, differing only in prompts. Currently, the boundaries between various big data tasks are unclear, and many tasks can be converted into generative tasks. The popularity of ChatGPT has even given rise to a new role, "prompt engineer," who specializes in crafting effective prompts to leverage ChatGPT for task completion. In the future, large models can liberate and enhance productivity, handling repetitive tasks and freeing humans for creative work, thus advancing technological progress.
E.3 Comparison Between Industrial and Research Institutions in Large Model Development
Industries aim to generate profits, driving them to develop market-appropriate large models with high enthusiasm. They also have vast user bases that provide data and accelerate training. However, industries face ethical and legal constraints, such as patents and privacy issues. Research institutions, lacking immediate application needs or access to first-hand data, face disadvantages in a data-driven era. Nonetheless, they pursue innovation without profit motives, experimenting in numerous directions, where even a one-in-a-million success marks a significant breakthrough. Furthermore, research institutions frequently refresh their teams, maintaining high efficiency in research.
Appendix F Potential Future Work
F.1 Dataset Extension
Collecting data from multiple sources can help mitigate the bias faced in the dataset.
F.2 Intelligent Triage
The All-in-one LLMs can solve the triage in a more general way. Researchers may pay more attention to adding more background information, such as multi-modality methods. Double checking will ground truth after the decision is also important due to the high risk of medical problems.
F.3 Medical Consultant
Medical consultation systems have a promising future as AI technology advances, with human resources and potentially unlimited computing resources that can scale with demand. With the development of current LLMs, constructing a consultation system become easier. Researchers can shift their focus to the alignment between LLM output and human preference (legal issues, correct knowledge, cost trade-off). The RLHF may contribute to the alignment.
F.4 Further Application
The power of LLMs may empower more complicated tasks. As a result, researchers may not need to formulate real tasks into fixed mathematical tasks, which will enable more flexible and realistic solutions. Researchers may extend the usage of our datasets based on their own need.

### v2

@@
-LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation
+Building a Chinese Medical Dialogue System: Integrating Large-scale Corpora and Novel Models
 Abstract
@@
 Second, existing methods lack medical knowledge and struggle to accurately understand professional terms and expressions in patient-doctor consultations.
-To overcome these obstacles, we construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC), comprising a Coarse-grained Triage dataset with 439,630 samples, a Fine-grained Diagnosis dataset with 199,600 samples, and a Medical Consultation dataset with 472,418 items, thereby addressing the data shortage in this field.
-Moreover, we further propose a novel triage system that combines BERT-based supervised learning with prompt learning, as well as a GPT-based medical consultation model using reinforcement learning.
+To overcome these obstacles, we construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC) , thereby addressing the data shortage in this field.
+Moreover, we further propose a novel triage system that combines BERT-based supervised learning with prompt learning, as well as a GPT-based medical consultation model.
 To enhance domain knowledge acquisition, we pre-trained PLMs using our self-constructed background corpus.
 Experimental results on the LCMDC demonstrate the efficacy of our proposed systems.
-Our corpora and codes are available at: Corpora and Codes
-LCMDC: Large-scale Chinese Medical Dialogue Corpora for Automatic Triage and Medical Consultation
-Xinyuan Wang  and Haozhou Li  and Dingfang Zheng  and Qinke Peng Xi’an Jiaotong University {wxy0713, lihaozhou1126, zdf0921}@stu.xjtu.edu.cn qkpeng@mail.xjtu.edu.cn
-1 Introduction
-The outbreak and spread of COVID-19 have placed immense pressure on global healthcare systems, exposing numerous shortcomings in traditional healthcare services Sun et al.
-(2021).
-In response to the urgent need for pandemic control, online medical consultations and tele-medicine have become crucial for maintaining healthcare services Jnr (2020).
+Our corpora and codes are available at: Corpora and Codes.
+Index Terms:
+Pre-trained Language Models, Medical Triage, Medical Consultation, Domain Knowledge Acquisition
+I Introduction
+The outbreak and spread of COVID-19 have placed immense pressure on global healthcare systems, exposing numerous shortcomings in traditional healthcare services [1].
+In response to the urgent need for pandemic control, online medical consultations and telemedicine have become crucial for maintaining healthcare services [2].
 Between July and December 2022, 58.5% of U.S.
-adults used the Internet to seek health or medical information, 41.5% communicated with a doctor or doctor’s office, and 46.1% checked medical test resultsWang and Cohen (2023).
+adults used the Internet to seek health or medical information, 41.5% communicated with a doctor or doctor’s office, and 46.1% checked medical test results[3].
 According to the China Internet Network Information Center (CNNIC), as of December 2023, the number of Internet medical users in China reached 414 million.
 This trend reflects the growing reliance on digital platforms for healthcare.
-The rise of online healthcare platforms has made medical triage and intelligent Q&A key research areas, helping reduce the burden on healthcare systems by providing remote medical advice and reducing unnecessary hospital visits.
-These technologies enhance resource efficiency, minimize cross-infection risks, and improve access to healthcare services, especially in remote areas.
+The rise of online healthcare platforms has made medical triage and intelligent Q&A key research areas, helping reduce the burden on healthcare systems, providing remote medical advice, and reducing unnecessary hospital visits.
+These technologies enhance resource efficiency, minimize cross-infection risks, and improve access to healthcare services.
 However, most existing studies face two main challenges.
-Firstly, due to data privacy concerns, obtaining large-scale public medical triage and consultation datasets is difficult.
-Although some studies have collected corpora for automatic diagnosis Wei et al.
-(2018); Lin et al.
-(2019); Xu et al.
-(2019) and dialogue generation Liu et al.
-(2022), these datasets are limited to a few specific diseases and are small in size Zeng et al.
-(2020), making them inadequate for training intelligent medical diagnosis models based on pre-trained language models.
-Secondly, patient-doctor consultations usually utilize distinct expressions and professional terms, making it difficult for existing methods that lack domain-specific knowledge to comprehend medically pertinent information Li et al.
-(2024).
-To address the limitations of existing datasets, we have constructed the Large-scale Chinese Medical Dialogue Corpora (LCMDC) by extracting patient-doctor conversations from the "Quick Doctor" website  [^1] https://www.120ask.com/.
+Firstly, due to data privacy concerns, obtaining complete large-scale public medical triage and consultation datasets is difficult [4].
+Although some studies have collected corpora for automatic diagnosis [5, 6, 7] and dialogue generation [8], these datasets are limited to a few specific diseases and are small in size [9], making them inadequate for training intelligent medical diagnosis models based on pre-trained language models.
+Secondly, patient-doctor consultations usually utilize distinct expressions and professional terms, making it difficult for existing methods that lack domain-specific knowledge to comprehend medically pertinent information [10].
+To address the limitations, we have built a Chinese medical dialogue system, including 1) the medical corpora, 2) the medical triage system, and 3) the medical consultation model.
+The Large-scale Chinese Medical Dialogue Corpora (LCMDC) extracts patient-doctor conversations from the "Quick Doctor" website  [^1] https://www.120ask.com/.
 Our work addresses two key needs in online healthcare: task-oriented medical triage and open-ended consultation systems.
-We developed a coarse-grained triage dataset with over 430,000 patient-doctor dialogues across 14 departments and a fine-grained diagnosis dataset with 200,000 entries covering 120 diseases.
-For medical consultation, we compiled over 470,000 question-answer pairs.
+We developed a coarse-grained triage dataset with over 430k patient-doctor dialogues across 14 departments and a fine-grained diagnosis dataset with 200k entries covering 120 diseases.
+For the consultation, we compiled over 470k question-answer pairs.
 Additionally, we collected disease information datasets to enrich Pre-trained Language Models (PLMs) with medical knowledge for pre-training.
-Nowadays, natural language processing techniques are increasingly effective in intelligent medical triage and dialogue generation tasksSun et al.
-(2017).
-Specifically, neural network-based text classification algorithms with attention mechanisms have been successfully applied to medical triage and diagnosisWei et al.
-(2018); Lin et al.
-(2019); Xu et al.
-(2019).
-However, existing studies often concentrate on a limited number of diseases and struggle with imbalanced label distributions in many-class medical diagnoses Zeng et al.
-(2020).
-Although some pre-trained language models, such as BERT and ERNIE, have been employed in automatic triage systems Wang et al.
-(2021); Wang and Feng (2022), they lack pre-training on medical data, hindering their ability to grasp domain-specific knowledge Li et al.
-(2024).
+Nowadays, natural language processing techniques are increasingly effective in intelligent medical triage and dialogue generation tasks[11, 12].
+Specifically, neural network-based text classification algorithms with attention mechanisms have been successfully applied to medical triage and diagnosis[5, 6, 7].
+However, existing studies often concentrate on a limited number of diseases and struggle with imbalanced label distributions in many-class medical diagnoses [9].
+Although some pre-trained language models, such as BERT and ERNIE, have been employed in automatic triage systems [13, 14], they lack pre-training on medical data, hindering their ability to grasp domain-specific knowledge [10].
 To overcome these limitations, we propose a novel triage system that combines PLM-based supervised learning with prompt learning methods.
-We initially pre-train BERT using a self-built medical background corpus, then employ medical BERT, BiLSTM, and Dendrite Network Liu and Wang (2021) to represent the input questions for classification in our Triage datasets.
+We initially pre-train BERT using a self-built medical background corpus, then employ medical BERT, BiLSTM, and Dendrite Network [15] to represent the input questions for classification in our Triage datasets.
 Additionally, we incorporate prompt learning to mitigate the many-label imbalance issue, enhancing prediction accuracy for rare diseases.
-In the medical text generation task, sequence-to-sequence models, including LSTM-based Seq2Seq methods Liu et al.
-(2022) and transformer-based PLMs like BERT-GPT Zeng et al.
-(2020), have been widely used, and achieved promising results.
+In the medical consultation model, sequence-to-sequence models, including LSTM-based Seq2Seq methods [8] and transformer-based PLMs like BERT-GPT [9], have been widely used, and achieved promising results.
 However, traditional LSTM models, which generate output solely based on input sequences, struggle with insufficient information utilization due to their gating mechanisms.
@@
 We also construct a domain knowledge graph and extract relevant content about each user’s question as supplementary input information to enhance the model.
-Finally, we train our model using our Medical Consultation Dataset and fine-tune it with a reinforcement learning method to simulate human doctors in providing responses.
+Finally, we train our model using our Medical Consultation Dataset.
 The contributions of this paper are three-fold:
@@
   • Proposed an information fusion-based classification algorithm and a prompt learning classification algorithm to improve intelligent triage accuracy, especially for rare diseases, achieving a 5% improvement in accuracy.
-  • Developed a dialogue model construction framework including knowledge incorporation and input supplementation, improving performance across multiple evaluation metrics.
-2 Chinese Medical Corpus
+  • Developed a medical consultation model including knowledge incorporation and input supplementation, improving performance across multiple evaluation metrics.
+II Related Work
+The development of AI-driven medical dialogue systems benefits from advancements in biomedical research and feature engineering techniques.
+Studies on bioinformatics and medicine [16, 17, 18, 19, 20, 21] provide valuable domain knowledge that can enhance medical AI applications.
+Additionally, effective feature selection and transformation techniques are crucial for optimizing model performance.
+Feature selection methods [22, 23, 24, 25] help identify the most informative clinical attributes, while feature transformation approaches [26, 27, 28, 29, 30] improve data representation, enabling more accurate and robust medical triage and consulting systems.
+Furthermore, large language models (LLMs) have significantly advanced research in medical AI by improving the understanding and processing of complex biomedical texts [31, 32, 33].
+Pre-trained models with domain-specific fine-tuning enable more effective medical dialogue generation, question answering, and clinical decision support.
+II-A Dataset
+CBLUE [34] is a comprehensive Chinese biomedical language understanding benchmark designed to advance research in biomedical NLP for the Chinese language.
+MedDG [8] is a large-scale dataset aimed at developing conversational agents for primary clinical advice, especially valuable during the COVID-19 pandemic.
+MedDialog [9] is a comprehensive dataset for training and evaluating medical dialogue systems, enhancing remote healthcare services.
+MedQuAD [35] is a significant resource for developing AI models in medical question answering, providing a robust foundation for understanding and responding to medical inquiries.
+There are some other datasets like event recognition [36].
+II-B Triage System
+IntelTriage [37] is an intelligent triage system for Emergency Departments in Greece, dynamically prioritizing patients and monitoring their vital signs using wearable biosensors.
+The Babylon AI-powered Triage and Diagnostic System [38] demonstrated diagnostic accuracy comparable to human doctors and provided safer triage recommendations.
+MESTRIMAN [39] is an expert system designed to support medical assistance during catastrophes by coordinating resources and enhancing disaster management.
+The SADE model [10] addresses challenges in medical triage and diagnosis by using dual encoders to represent patient consultations and doctor diagnoses.
+II-C Question Answer System
+This research [40] automates the generation of qualified medical answers using an end-to-end RNN model trained on data from online health services.
+Another system [41] integrates medical knowledge, knowledge graphs, and QA systems, using crawler technology and rule-based algorithms for classification and querying.
+An end-to-end character-level multi-scale CNN framework [42] is introduced for Chinese medical question-answer matching.
+Med-PaLM 2 [43] leverages large language model improvements and medical domain fine-tuning to enhance performance on medical QA datasets.
+III Chinese Medical Corpus
 To address the scarcity of large-scale corpora and the constraints of current datasets, we collected millions of doctor-patient conversations to construct the Large-scale Chinese Medical Dialogue Corpora (LCMDC).
 This collection is designed to enhance medical triage and dialogue generation and includes three components: 1) a Coarse-grained Triage dataset, 2) a Fine-grained Diagnosis dataset, and 3) a Medical Consultation dataset.
-2.1 Data Collection
+III-A Data Collection
 To make the model more adaptable to conversational and ambiguous medical descriptions, we claw the original data from the online medical consultation platform named "Quick Doctor".
-Reasons are shown in Appendix B.1.
+Here are several reasons:
+  • Apart from sensitive information, the patient consultations and doctor responses are public, reducing the difficulty of data collection.
+  • The portal boasts an extensive dataset, containing over 100 million entries, which adequately meets our requirements.
+  • The portal provides comprehensive information, including user characteristics like gender and age, and detailed classifications of medical departments across four levels, facilitating robust data statistics and medical department recommendations.
 (a) Age Distribution
@@
 The results show that female users significantly outnumber male users.
-Predominantly, young individuals aged 10-40, who are more likely to incorporate the Internet into their daily routines, made the inquiries.
-2.2 Intelligent Triage Dataset
+Predominantly, young individuals aged 10-30, who are more likely to incorporate the Internet into their daily routines, made the inquiries.
+III-B Intelligent Triage Dataset
 For automatic medical triage and diagnosis, we construct two datasets of different granularities for experimental analysis.
@@
 We organize the original corpus into two datasets after data preprocessing: 1) a Coarse-grained Triage dataset with primary department labels for higher accuracy in quickly identifying potential conditions and directing patients to the appropriate department, and 2) a Fine-grained Diagnosis dataset with tertiary department labels for more precise guidance, thereby helping users access detailed medical knowledge and improving query efficiency.
-The Coarse-grained dataset has over 430,000 entries with 14 categories, while the Fine-grained dataset has over 190,000 entries with 120 categories, as shown in Table 1.
-Based on these datasets, we further propose an intelligent triage algorithm based on supervised learning and prompt-based learning and then evaluated its effectiveness.
-Table 1: Statistic of Intelligent Triage Datasets
+The Coarse-grained dataset has over 430k samples with 14 categories, while the Fine-grained dataset has over 190k samples with 120 categories, as shown in Table I.
+Based on these datasets, we further propose an intelligent triage algorithm based on supervised learning and prompt-based learning and then evaluate its effectiveness.
+TABLE I: Statistic of Intelligent Triage Datasets
 Dataset Name Coarse-grained Fine-grained
@@
 Number of Categories 14 120
-2.3 Medical Consultation Dataset
+III-C Medical Consultation Dataset
 For the medical dialogue generation, we extract the Question-Answer sample from the original corpus to create the Medical Consultation Dataset, which can offer several advantages: 1) Major medical Q&A portals have a large user base with comprehensive questions, providing a wide range of input data for the model to learn from; 2) High-quality doctors registered on these sites provide accurate responses, providing high-quality domain knowledge for training a professional model; 3) The large data volume allows extensive training to improve model performance from various perspectives.
-Table 2: Statistic of Medical Consulting Dataset
+TABLE II: Statistic of Medical Consulting Dataset
 Dataset Item Quantity
@@
 After data preprocessing, the Medical Consultation Dataset contains a total of 472,418 samples, with an average question length of 57.86 characters and an average response length of 92.57 characters.
-The statistical results of the dataset are presented in Table 2.
-3 Intelligent Triage System
+The statistical results are presented in Table II.
+IV Intelligent Triage System
 We propose two text classification-based methods to evaluate intelligent triage datasets, using coarse-grained and fine-grained data to provide varying levels of consultation suggestions.
@@
 For categories with limited data, a prompt-based text classification method leverages pre-trained language models’ knowledge, enhancing performance on small datasets.
-3.1 Problem Definition
+IV-A Problem Definition
 The intelligent medical triage task is modeled as a text classification problem, processing user queries to determine the appropriate department for consultation.
 $p\left\{{c_{1},~{}c_{2},\ldots,c_{k_{l}}}\right\}=TextCls\left({t_{1},t_{2},\ldots,t_{k_{w}}}\right),$ (1)
-where $p\left\{{c_{1},~{}c_{2},\ldots,c_{k_{l}}}\right\}$ means the probability on each label.
-3.2 Supervised Learning Triage System
+IV-B Supervised Learning Triage System
 First, we propose a traditional supervised learning-based text classification method.
@@
 $loss=-{\sum\limits_{i=1}^{k_{l}}{c_{i}{\mathit{\log}\left(p_{i}\right)}}}$ (10)
-3.3 Prompt Learning Triage System
+IV-C Prompt Learning Triage System
 Both granularity-level classification datasets suffer from data imbalance issues.
@@
 This accelerates the training process and is suitable for categories with fewer training samples.
-4 Medical Consultation System
+V Medical Consultation System
 For the medical consultation dataset, we developed a Chinese medical dialogue system using a generative language model to simulate interactions with doctors.
 The system incorporates medical knowledge and supplements inputs, with user questions as input text and doctor’s responses as output text for training.
-4.1 Problem Definition
+V-A Problem Definition
 We mathematically model building a consultation system as a text-generation task.
@@
 $=TextGen\left(q_{1},q_{2},\ldots,q_{M},i_{1},i_{2},\ldots,i_{L}\right)$
-4.2 Overview
+V-B Framework Overview
 Based on the GPT-2 model, we re-train the model using the collected medical encyclopedia text data to internalize knowledge, making it more suitable for tasks in the medical domain.
 Next, we construct a medical knowledge graph and use a search-and-match approach to find relevant knowledge to each question, providing external information as supplementary input text.
-Then, we train the model using dialogue data and finally fine-tune it using reinforcement learning.
+Then, we train the model using dialogue data.
 According to the above process, we construct the domain-specific medical consultation system, as shown in Figure 4.
 Figure 4: The Medical Dialogue System Framework.
-4.3 Knowledge Injection Answer Generation
+V-C Knowledge Injection Answer Generation
 Knowledge injection allows the model to have rich prior knowledge, which involves re-training the large language model using the collected basic knowledge text data.
@@
 In the re-training stage, we view a sentence as a token sequence: $Sentence=\left({{token}_{1},{token}_{2},\ldots,{token}_{N}}\right)$. The objective is to maximize the following objective function:
-$L\left(Sentence\right)=\sum\limits_{n=K}^{N}p\left({token}_{n}\,\middle|\,{token}_{n-K},\ldots,{token}_{n-1};\theta\right)$ (12)
+$\scriptsize L\left(Sentence\right)=\sum\limits_{n=K}^{N}p\left({token}_{n}\,\middle|\,{token}_{n-K},\ldots,{token}_{n-1};\theta\right)$ (12)
 where $\theta$ represents the parameters in the GPT-2 model that need to be fine-tuned, and encompass knowledge.
@@
 This corresponds to the fine-tuning phase of the pre-trained language model, enabling the model to simulate human doctors in responding.
-5 Experiments
+VI Experiments
 We build the intelligent triage system and medical consultation system and test them on the collected datasets.
-5.1 Experiment Setup
-The experiments for the triage and consultation systems, including datasets, baselines, evaluation metrics, hyper-parameters, and environment settings, are outlined in detail in the Appendix C.
-5.2 Supervised Triage System Experiment
-5.2.1 Overall Performance
+VI-A Experiment Setup
+VI-A1 Datasets
+We used the datasets shown in Section III, including the intelligent triage and consultant datasets.
+VI-A2 Baselines
+We employ the mechanism learning baselines, neural network baselines, and pre-trained language models, shown in Table III and Table V.
+VI-A3 Evaluation Metrics
+For the triage system, we utilize accuracy, precision, recall, and F-1 score.
+For the consultation system, we employ n-gram overlap, distance-based, vector similarity, and sentence semantic similarity measurements.
+VI-A4 Hyper Parameters
+Continuous pre-training: Max length 128, mask 15% tokens in MLM.
+Supervised training: Embedding dim 768, LSTM hidden size 1024, 50 epochs.
+PTM lr 5e-5, others 2e-4.
+Prompt learning: 20 epochs, lr 2e-5.
+Consultation system: Base model GPT-2, token embedding dim 768, vocab length 50257.
+Pre-train 10 epochs with background knowledge, fine-tune 20 epochs, lr 2.6e-5.
+VI-A5 Experimental Environment
+All experiments were conducted on Ubuntu 22.04.3 LTS OS, with the framework of Python 3.11.5 and PyTorch 2.0.1.
+All data are computed on 4 NVIDIA GeForce RTX 2080 Ti GPUs, each featuring 12 GB of memory with CUDA version 12.2.
+VI-B Supervised Triage System Experiment
+VI-B1 Overall Performance
 For the supervised learning classification method, we employ different baselines including machine learning methods, neural network methods, and pre-trained models.
-The results of different-grained datasets are shown in Table 3
-Table 3: Experimental Results on Different Granularity Datasets
+The results of different-grained datasets are shown in Table III.
+TABLE III: Experimental Results on Different Granularity Datasets
 Coarse-grained Dataset Results Fine-grained Dataset Results
@@
 Decision Tree 61.24% 48.86% 52.47% 45.71% 59.68% 45.61% 45.49% 45.74%
-Neural Network TextCNN 77.07% 54.36% 56.83% 52.10% 76.38% 61.29% 66.92% 56.54%
-RCNN 71.56% 51.40% 53.76% 49.24% 73.84% 59.86% 61.10% 58.67%
+Neural Network TextCNN [44] 77.07% 54.36% 56.83% 52.10% 76.38% 61.29% 66.92% 56.54%
+RCNN [45] 71.56% 51.40% 53.76% 49.24% 73.84% 59.86% 61.10% 58.67%
 LSTM 73.38% 53.02% 55.37% 50.86% 71.59% 59.80% 66.91% 54.08%
 Self Attention 64.52% 52.17% 54.22% 50.27% 68.08% 60.17% 64.90% 56.08%
-Capsule 62.70% 51.47% 53.01% 50.01% 65.98% 52.23% 57.88% 47.59%
+Capsule [46] 62.70% 51.47% 53.01% 50.01% 65.98% 52.23% 57.88% 47.59%
 Pre-trained Model BERT 79.60% 59.92% 62.51% 58.79% 80.53% 65.65% 70.64% 65.24%
-RoBERTa 77.59% 55.69% 60.23% 53.98% 77.31% 55.74% 61.37% 53.46%
-MacBERT 80.17% 59.68% 66.56% 59.84% 80.73% 65.23% 71.13% 64.88%
+RoBERTa [47] 77.59% 55.69% 60.23% 53.98% 77.31% 55.74% 61.37% 53.46%
+MacBERT [48] 80.17% 59.68% 66.56% 59.84% 80.73% 65.23% 71.13% 64.88%
 Experimental Algorithm No PT Benchmark 81.34% 63.20% 65.05% 62.76% 82.47% 72.01% 75.60% 69.78%
@@
 This results in better classification outcomes on the fine-grained dataset compared to the coarse-grained dataset.
-Ablation study results are presented in Appendix D.1
-5.2.2 Results of Domain Information Injection
+VI-B2 Ablation study
+When any part of the proposed classification algorithm is removed or replaced with another structure, the classification accuracy of the new model decreases, as shown in Figure 5.
+This indirectly highlights the importance of each component of the proposed model.
+(a) Coarse-grained
+(b) Fine-grained
+Figure 5: Ablation Results of Supervised Classification Method.
+VI-B3 Results of Domain Information Injection
 To investigate the differences before and after domain-specific re-training of different pre-trained models, we designed comparative experiments to evaluate model performance under different conditions.
 We used the original BERT and MacBERT models, as well as the proposed models utilizing these pre-trained models, comparing their performance before and after re-training.
-The results are shown in Figure 5
+The results are shown in Figure 6.
 (a) Coarse-grained
@@
 Figure 6: Results of Domain Information Injection.
-With the same pre-trained models and classification network structures, re-training the language model can improve the accuracy of the classification model.
-This is because the re-training data is related to Chinese medical information, allowing the language model to learn the intrinsic knowledge of medical texts, resulting in better performance on tasks involving medical texts.
+With the same pre-trained models and classification network structures, re-training the language model can improve the classification accuracy.
+This is because the re-training data is related to Chinese medical information, allowing the language model to learn the intrinsic knowledge of medical texts, resulting in better performance of medical texts.
 From this set of comparative data, it can be seen that using re-trained language models significantly improves accuracy and F1 scores.
@@
 In this experiment, the language model re-trained with medical knowledge and Q&A text data demonstrates better representation capability for texts that integrate colloquial expressions and medical knowledge.
-5.2.3 Parameters Study
-We also investigated the impact of different numbers of dendritic neural network layers and LSTM layers on classification results to select the optimal algorithm model structure.
+VI-B4 Parameters Study
+We also investigated the impact of different numbers of layers of dendritic neural network and LSTM layers on classification results to select the optimal algorithm model structure.
 (a) DD layer
@@
 Figure 7: Results of DD and LSTM layer Numbers.
-According to the results in Figure 6, the dendritic neural network achieves the best classification performance with 3 layers, and the LSTM network achieves optimal performance with 2 layers.
-5.3 Prompting Triage System Experiment
+According to Figure 7, the dendritic neural network achieves the best classification performance with 3 layers, and the LSTM network achieves optimal performance with 2 layers.
+VI-C Prompting Triage System Experiment
 Prompt learning quickly completes tasks using pre-trained knowledge, even with limited data.
 To leverage this advantage, we created small sample datasets by removing high-data categories from the coarse-grained and fine-grained datasets, resulting in 73,604 entries across nine categories and 24,940 entries across 83 categories, respectively.
-Table 4: Prompt Learning Classification Results on Different Granularity Datasets
+TABLE IV: Prompt Learning Classification Results
 Dataset Method Model
@@
 Small Sample 79.28% 80.11% 77.31%
-Table 4 shows that the BERT model re-trained with collected data performs better in prompt learning-based classification, surpassing supervised learning on small datasets.
-However, prompt learning has drawbacks: it requires more training time due to predicting multiple token positions, and its performance can be slightly worse than supervised learning.
+Table IV shows that the BERT model re-trained with collected data performs better in prompt learning-based classification, surpassing supervised learning on small datasets.
+However, prompt learning has drawbacks: it requires more time due to predicting multiple tokens, and its performance can be slightly worse than supervised learning.
 This is due to the need to correctly predict all token positions and handle varying label lengths with padding symbols, which can interfere with pre-trained knowledge and reduce accuracy.
 Additionally, prompt learning is less effective with large datasets.
-5.4 Question Answer System Experiment
-5.4.1 Overall Performance
+VI-D Medical Consultation System Experiment
+VI-D1 Overall Performance
 First, we compared the proposed framework’s text generation model with a deep learning-based sequence-to-sequence (seq2seq) model.
@@
 In this study, we utilized the general Chinese BART model to build a dialogue system and trained it using medical dialogue data.
-Table 5: Comparison of Text Generation Results
+TABLE V: Comparison of Text Generation Results
 Evaluation Metric Seq2Seq Model BART Our Model
@@
 Weight R $\uparrow$ 0.1321 0.1998 0.2102
-BLEU-1 $\uparrow$ 0.0922 0.1214 0.1211
-CHRF $\uparrow$ 0.0374 0.0769 0.0781
-GLEU $\uparrow$ 0.0448 0.0977 0.1029
-NIST $\uparrow$ 0.5398 1.3932 1.4049
-RIBES $\uparrow$ 0.0637 0.0827 0.0903
-TER $\downarrow$ 1.3519 1.0489 0.9997
-WMD $\uparrow$ 0.5168 0.6192 0.6461
-BERT F1 $\uparrow$ 0.568 0.646 0.648
+BLEU-1 [49] $\uparrow$ 0.0922 0.1214 0.1211
+CHRF [50] $\uparrow$ 0.0374 0.0769 0.0781
+GLEU [51] $\uparrow$ 0.0448 0.0977 0.1029
+NIST [52] $\uparrow$ 0.5398 1.3932 1.4049
+RIBES [53] $\uparrow$ 0.0637 0.0827 0.0903
+TER [54] $\downarrow$ 1.3519 1.0489 0.9997
+WMD [55] $\uparrow$ 0.5168 0.6192 0.6461
+BERT F1 [56] $\uparrow$ 0.568 0.646 0.648
 BERT P $\uparrow$ 0.576 0.649 0.652
@@
 Web Bert Similarity $\uparrow$ 2.2324 2.2858 2.2960
-According to Table 5, the proposed model outperforms the traditional deep learning sequence-to-sequence (seq2seq) algorithm across almost all evaluation metrics.
+According to Table V, the proposed model outperforms the traditional deep learning sequence-to-sequence (seq2seq) algorithm across almost all evaluation metrics.
 TER (Translation Edit Rate) measures the number of edits required to change a generated sentence to match the reference sentence, indicating the proximity between the generated and real texts.
@@
 However, these characters do not add actual meaningful content, as reflected in the other metrics.
-Ablation study results are presented in Appendix D.2
-6 Discussion and Future Work
-We discuss the biases in online medical data, AI adoption in healthcare, and the development of large language models in different sectors. For details, please refer to Appendix E.
-Looking ahead, future work focuses on expanding datasets to reduce bias, enhancing intelligent triage systems with background knowledge and multi-modal methods, and refining medical consultation systems through alignment between LLM outputs and human preferences.
-The potential for further applications of LLMs is vast, enabling more flexible and realistic solutions in various fields.
-For details, please refer to the Appendix F.
-7 Conclusion
-We collected and organized Chinese medical data, including datasets for automated medical triage and consultation, forming the foundation of our work.
-Based on these datasets, we developed two systems: an intelligent triage system using text classification and a medical consultation system using a generative language model.
-Experimental results and ablation studies validated the effectiveness of both systems, demonstrating improved performance in medical diagnosis and consultation tasks.
-8 Limitation
-Dataset Biases:
-  • Lack of Emergency Data: The dataset lacks emergency and severe cases, limiting the model’s ability to generalize to critical or urgent medical situations.
-  • Imbalanced Age Distribution: The dataset is biased towards younger users, with insufficient data on elderly populations, reducing the model’s applicability to older age groups.
-  • Issues Due to Anonymity: The anonymity of online consultations often leads to more private or informal questions, which may not accurately reflect typical clinical scenarios, affecting the model’s real-world performance.
-Intelligent Triage System:
-  • Lack of Comparison with Human Performance: Current classification methods lack sufficient comparison with professional doctors’ diagnostic results, making assessing their reliability in practical medical applications difficult.
-  • Lack of Multi-modal Research: Existing classification methods are primarily text-based and do not incorporate multi-modal data such as images or audio, limiting the understanding of comprehensive patient information.
-Medical Consultation System:
-  • No Multi-turn Dialogues: Generation methods are often limited to single-turn dialogues, lacking the ability to handle contextual information across multiple rounds of interaction, which is crucial in real consultations.
-  • No Role-playing or Simulation: There is a lack of role-playing or simulation of real medical scenarios in the generated dialogues, resulting in responses that lack authenticity and practical relevance to clinical settings.
-9 Ethical Issues
-Our study does not present any ethical issues that would require formal ethics approval, as it does not involve human subjects, sensitive personal data, or any other ethical considerations that would necessitate formal ethics oversight.
-First, the anonymized patient-doctor dialogues were legally and compliantly obtained from publicly available sources, specifically from an online medical consultation platform.
-Second, users have agreed to the platform’s terms of service, which include being informed that these anonymized consultation dialogues may be made public, so the data has been made available to the public with the informed consent of the users.
-Appendix A Related Work
-A.1 Dataset
-CBLUE Zhang et al.
-(2022) is a comprehensive Chinese biomedical language understanding benchmark designed to advance research in biomedical NLP for the Chinese language.
-MedDG Liu et al.
-(2020) is a large-scale dataset aimed at developing conversational agents for primary clinical advice, especially valuable during the COVID-19 pandemic.
-MedDialog Zeng et al.
-(2020) is a comprehensive dataset for training and evaluating medical dialogue systems, enhancing remote healthcare services.
-MedQuAD Ben Abacha and Demner-Fushman (2019) is a significant resource for developing AI models in medical question answering, providing a robust foundation for understanding and responding to medical inquiries.
-There are some other datasets like event recognition Deng et al.
-(2022).
-A.2 Triage System
-IntelTriage Billis et al.
-(2019) is an intelligent triage system for Emergency Departments in Greece, dynamically prioritizing patients and monitoring their vital signs using wearable biosensors.
-The Babylon AI-powered Triage and Diagnostic System Razzaki et al.
-(2018) demonstrated diagnostic accuracy comparable to human doctors and provided generally safer triage recommendations.
-MESTRIMAN Sierra et al.
-(1993) is an expert system designed to support medical assistance during catastrophes by coordinating resources and enhancing disaster management.
-The SADE model Li et al.
-(2024) addresses challenges in medical triage and diagnosis by using dual encoders to represent patient consultations and doctor diagnoses.
-A.3 Question Answer System
-This research Abdallah et al.
-(2020) automates the generation of qualified medical answers using an end-to-end RNN and encoder-decoder model trained on data from online health services.
-Another system Jiang et al.
-(2021) integrates medical knowledge, knowledge graphs, and QA systems, using crawler technology and rule-based algorithms for classification and querying.
-An end-to-end character-level multi-scale CNN framework Zhang et al.
-(2017) is introduced for Chinese medical question-answer matching.
-Med-PaLM 2 Singhal et al.
-(2023) leverages large language model improvements and medical domain finetuning to enhance performance on medical QA datasets.
-Appendix B Dataset Details
-B.1 Reasons for Data Source
-Here are several reasons for choosing "Quick Doctor" website:
-  • Apart from sensitive information, the patient consultations and doctor responses are public, reducing the difficulty of data collection.
-  • The portal boasts an extensive dataset, containing over 100 million entries, which adequately meets our requirements.
-  • The portal provides comprehensive information, including user characteristics like gender and age, and detailed classifications of medical departments across four levels, facilitating robust data statistics and medical department recommendations.
-Appendix C Experiment Setup Details
-C.1 Datasets
-In Section 2, we collect medical corpus and formulate 2 datasets for the triage system and consultation system respectively.
-In the intelligent triage dataset, the inputs are the users’ questions, and the outputs are two different grained labels, indicating high-level and detailed triage.
-C.2 Baselines
-For the triage system, we employ SVM Cortes and Vapnik (1995) and decision tree Quinlan (1986) as the mechanism learning baselines, while TextCNN Kim (2014), RCNN Lai et al.
-(2015), LSTM Hochreiter and Schmidhuber (1997), Self-Attention Vaswani et al.
-(2017), and Capsule Network Sabour et al.
-(2017) are used as neural network baselines.
-Then We also utilize pre-trained language models including BERT, RoBERTa Liu et al.
-(2019), and MacBERT Cui et al.
-(2020).
-For the consultation system, we employ the Seq2Seq method and BART Lewis et al.
-(2020) as baselines.
-C.3 Evaluation Metrics
-For the triage system, we utilize classification metrics, including accuracy, precision, recall, and F-1 score.
-For the consultation system, we employ various evaluations, including n-gram overlap measures such as BLEU Papineni et al.
-(2002), NIST Doddington (2002), RIBES Isozaki et al.
-(2010), CHRF Popović (2015), and GLEU Mutton et al.
-(2007); distance-based measures like TER Snover et al.
-(2006); vector similarity measures including WMD Kusner et al.
-(2015); n-gram diversity measures such as Self BLEU; sentence semantic similarity measures including BERT score Zhang et al.
-(2019) and Web Bert Similarity; and other metrics such as Lexical Diversity, and KL Divergence.
-C.4 Hyper Parameters
-For continuous pre-training of BERT and MacBERT, we set the max input length as 128 and mask 15% of tokens in the MLM task.
-For the supervised training of the triage system, we set the embedding dimension to 768, and the LSTM hidden size to 1024.
-We train 50 epochs.
-The learning rate for the pre-trained language model is 5e-5, while that for other parts of the whole framework is 2e-4.
-For the prompt learning method, we train 20 epochs with a learning rate of 2e-5.
-For the consultation system, we use GPT-2 as the base model.
-We set the token embedding dimension to 768, and the vocabulary length of natural language tokens to 50257.
-In the pre-training stage, we trained 10 epochs using background knowledge.
-In the fine-tuning stage, we trained 20 epochs using a learning rate of 2.6e-5.
-C.5 Experimental Environment
-All experiments were conducted on Ubuntu 22.04.3 LTS OS, with the framework of Python 3.11.5 and PyTorch 2.0.1.
-All data are computed on 4 NVIDIA GeForce RTX 2080 Ti GPUs, each featuring 12 GB of memory with CUDA version 12.2.
-Appendix D Ablation Study Details
-D.1 Supervised Triage System Experiment
-To verify the necessity of each component in the supervised triage system , we conducted ablation experiments by removing each component.
-(a) Coarse-grained
-(b) Fine-grained
-Figure 7: Ablation Results of Supervised Classification Method.
-When any part of the proposed classification algorithm is removed or replaced with another structure, the classification accuracy of the new model decreases, as shown in Figure 7.
-This indirectly highlights the importance of each component of the proposed model.
-D.2 Medical Consultation System Experiment
+VI-D2 Ablation study
 Next, we conducted ablation experiments to determine the importance of each component of the medical consultation system.
 Three comparative models were set up: a model without knowledge internalization, a model without input supplementation, and a model without both components.
-Table 6: Results of Text Generation Ablation Experiments
+TABLE VI: Results of Text Generation Ablation Experiments
 Evaluation Metric Ours w/o Knowledge w/o Input Neither
@@
 Web Bert Similarity $\uparrow$ 2.2960 2.2855 2.2819 2.2633
-The results in Table 6 demonstrate that each step in the proposed framework is crucial.
+The results in Table VI demonstrate that each step in the proposed framework is crucial.
 Removing any step results in a decline in the model’s performance across various evaluation metrics, highlighting the superiority and rationality of the proposed structure.
@@
 This process also contributes to the interpretability of the model construction process.
-Appendix E Discussion
-E.1 Data Issues
-During data collection, we found certain biases in the data gathered from online platforms.
-Firstly, acute and severe cases are not typically presented online, as patients in these situations usually go directly to the hospital.
-Secondly, the internet acts as a filter, with the online population differing from the general population.
-This results in a majority of younger users seeking online consultations, leading to a lack of data on elderly conditions.
-Lastly, due to the anonymity of the internet, the nature of questions asked online can differ from those asked offline, often including more private issues.
-Medical issues require maximum caution, and licensed doctors undergo extensive training to earn their qualifications due to the severe consequences of medical errors, which can endanger lives.
-Therefore, using current AI, which often functions as a black-box system, for diagnosis and treatment remains an ethical and legal grey area.
-At present, AI can only serve as a supplementary tool.
-Similar to autonomous driving, where accidents pose challenges in accountability, AI in healthcare must develop gradually.
-Automation systems must advance in a phased manner.
-E.2 Discussion on Current Large Models
-With the rise of ChatGPT, pre-trained language models seem to have deviated from their original purpose but have unified various tasks.
-Previously proposed task classifications are now outdated.
-Since the introduction of the T5 model, all-natural language processing tasks can be modeled as text generation tasks, differing only in prompts.
-Currently, the boundaries between various big data tasks are unclear, and many tasks can be converted into generative tasks.
-The popularity of ChatGPT has even given rise to a new role, "prompt engineer," who specializes in crafting effective prompts to leverage ChatGPT for task completion.
-In the future, large models can liberate and enhance productivity, handling repetitive tasks and freeing humans for creative work, thus advancing technological progress.
-E.3 Comparison Between Industrial and Research Institutions in Large Model Development
-Industries aim to generate profits, driving them to develop market-appropriate large models with high enthusiasm.
-They also have vast user bases that provide data and accelerate training.
-However, industries face ethical and legal constraints, such as patents and privacy issues.
-Research institutions, lacking immediate application needs or access to first-hand data, face disadvantages in a data-driven era.
-Nonetheless, they pursue innovation without profit motives, experimenting in numerous directions, where even a one-in-a-million success marks a significant breakthrough.
-Furthermore, research institutions frequently refresh their teams, maintaining high efficiency in research.
-Appendix F Potential Future Work
-F.1 Dataset Extension
-Collecting data from multiple sources can help mitigate the bias faced in the dataset.
-F.2 Intelligent Triage
-The All-in-one LLMs can solve the triage in a more general way.
-Researchers may pay more attention to adding more background information, such as multi-modality methods.
-Double checking will ground truth after the decision is also important due to the high risk of medical problems.
-F.3 Medical Consultant
-Medical consultation systems have a promising future as AI technology advances, with human resources and potentially unlimited computing resources that can scale with demand.
-With the development of current LLMs, constructing a consultation system become easier.
-Researchers can shift their focus to the alignment between LLM output and human preference (legal issues, correct knowledge, cost trade-off).
-The RLHF may contribute to the alignment.
-F.4 Further Application
-The power of LLMs may empower more complicated tasks.
-As a result, researchers may not need to formulate real tasks into fixed mathematical tasks, which will enable more flexible and realistic solutions.
-Researchers may extend the usage of our datasets based on their own need.
+VII Conclusion
+We collected and organized Chinese medical data, including datasets for automated medical triage and consultation, forming the foundation of our work.
+Based on these datasets, we developed two systems: an intelligent triage system using text classification and a medical consultation system using a generative language model.
+Experimental results validated the effectiveness of both systems, demonstrating improved performance in medical diagnosis and consultation tasks.
+Looking ahead, future work focuses on expanding datasets to reduce bias, enhancing intelligent triage systems with background knowledge and multi-modal methods, and refining medical consultation systems through alignment between LLM outputs and human preferences.
+At the same time, inspired by the field of recommendation systems [57, 58], further considering rich user information combined with the powerful generalization ability of LLM to achieve personalized triage is also a future direction.
+The potential for further applications of LLMs is vast, enabling more flexible and realistic solutions in various fields.